OpenStudy (anonymous):

i ^2 + i ^3 + i ^4 + i ^5 =

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

What did you get for i^2?

OpenStudy (anonymous):

mean i squared 2

OpenStudy (anonymous):

Right, but do you know how to figure out what i-squared is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i time i = squared i^2

OpenStudy (anonymous):

i is defined to be \[i=\sqrt{-1}\]so \[i^2=(\sqrt{1})^2=-1\] Does that make sense so far?

OpenStudy (anonymous):

i^n = i^(n mod4)

OpenStudy (anonymous):

Not sure if he knows modular arithmetic...

OpenStudy (anonymous):

For u, then:-)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yeah I remember - thanks!

OpenStudy (anonymous):

-1 + -i + 1 + i = 0

OpenStudy (anonymous):

Okay, so \[i^3=(i^2)i=(-1)i=-i\]

OpenStudy (anonymous):

Yes, Alchemista, we know, we are showing him how...

OpenStudy (anonymous):

yes 4^4 =1?

OpenStudy (anonymous):

i^4=1

OpenStudy (anonymous):

Yes - i^4= 1

OpenStudy (anonymous):

Yep, so now you just add them.

OpenStudy (anonymous):

i^5=-1

OpenStudy (anonymous):

i^2i^2i = (-1)(-1)i=i

OpenStudy (anonymous):

\[i^5=i^4*i=1*i=i\]

OpenStudy (anonymous):

That works, too.

OpenStudy (anonymous):

yes so its result is 0

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

No worries.